OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ Numerical Answers Expected! help please

OpenStudy (anonymous):

@strawberryswing

OpenStudy (solomonzelman):

well. \(\large\color{black}{ f_3=f_1+f_2+f_{3-1} }\) \(\large\color{black}{ f_3=2+4+f_{2} }\) \(\large\color{black}{ f_3=6+4 }\) \(\large\color{black}{ f_3=10 }\)

OpenStudy (solomonzelman):

then, \(\large\color{black}{ f_4=f_1+f_2+f_{4+1} ~~~\Rightarrow~~f_4=2+4+f_{3} ~~~\Rightarrow~~f_4=6+10=16 }\)

OpenStudy (anonymous):

ok thanx

OpenStudy (solomonzelman):

you see the pattern, right. (as the formula says it, but I want to make it visible, just simply) f(3) = 2+ 4 + 4 f(4) = 2+ 4 + 10 f(5) = 2+ 4 + 16

OpenStudy (solomonzelman):

you can re-write the recursive formula as: \(\large\color{black}{ f_{n}=6+f_{n-1} }\)

OpenStudy (solomonzelman):

in other words, \(\large\color{black}{ d=6 }\)

OpenStudy (anonymous):

ok thank u

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

btw, do you need a latex for sigma notation in text?

OpenStudy (solomonzelman):

`\(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ f(n)}\)` \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ f(n)}\)